OpenStudy (anonymous):

Please Help how to find intersection of x|3|2 y|3|4

OpenStudy (anonymous):

mid point or intersection because intersection requires 2 lines whereas i only see 2 coordinate points

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

\[x_{m} = \frac{x_2-x_1}{2}\] \[y_m = \frac{y_2-y_1}{2}\]

OpenStudy (anonymous):

how can i graph this